The government is mulling protecting the shoe industry. It estimates it needs to protect it for 6 years to the tune of $5 billion cost annually. Afterward, the shoe industry will return $200 million annually to the US society. Assume the repayment benefit lasts 100 years and the discount rate is 9%. Based on only economic reasoning and not political reasons, does it behoove the government to aid this industry? You need to explain lucidly what the analysis is. If you just compute numbers, it is not sufficient.
Answer #1

PV of cash outflows = Annuity*(1-1/(1+rate)^number of terms)/rate

= 5000000000*(1-1/(1+9%)^6)/9%

= 22429592951.15

PV of inflows at end of 6 years= Annuity*(1-1/(1+rate)^number of terms)/rate

= 200000000*(1-1/(1+9%)^100)/9%

= 2221820304.00

PV of inflows now = 2221820304/1.09^6 = $1,324,798,853.47

NPV = -22429592951.15+1324798853.47

= -21104794098

We see that the Net Present value added by this method is negative. Hence the project is not beneficial.